# Key Words.: stars: abundances – Local Group – galaxies: dwarf – galaxies: formation1 Introduction

### v1

@@ 4.2 Carbon
Figure 7 presents the carbon abundances measured in Fornax and Carina at [Fe/H] ¡ 2.5, compared to literature data. The four Carina stars analyzed in this paper all present very low carbon levels, similar to the upper limits derived in the Carina sample of Venn et al. (2012). The Carina [C/Fe] distribution is located at the lower edge of the Milky Way abundances, making Carina noticeably C-poor (see Fig. 4). However, we note that two CEMP-no stars have been identified in Carina, first in Susmitha et al. (2017), and more recently in Hansen et al. (2023).
In Fornax, both our stars and the one star previously studied by Tafelmeyer et al. (2010) have similar carbon-normal abundances. On average the Fornax stars are somewhat higher in C compared to Carina, especially when corrections for evolutionary status are taken into account.
We note that none of our six new stars are enhanced in carbon, while in the Milky Way the typical fraction of CEMP-no stars is at (Placco et al., 2014). This adds to previous results, showing that dSph galaxies are poor in CEMP-no stars relative to the Milky Way and the UFDs (Starkenburg et al., 2013; Skúladóttir et al., 2015, 2021; Skúladóttir et al., 2023a; Jablonka et al., 2015; Simon et al., 2015; Kirby et al., 2015; Hansen et al., 2018; Lucchesi et al., 2020), for a more detailed discussion on this see Section 5.
@@ 4.3 Light element: sodium
The LTE abundances of sodium are presented in Fig. 8. The Milky Way has a very large scatter in [Na/Fe] at low [Fe/H] ¡ , and our new data points fall within this range. While Fornax shows a near solar , all four stars in Carina have high . Since direct comparison between the Carina (X-Shooter) and Fornax (UVES) spectra is not meaningful, in Fig. 4 we plot the Na lines of the Carina stars together with an X-Shooter spectrum of a Na-poor star in Sculptor. This comparison clearly shows the Carina stars to be Na rich.
NLTE effects can play an important role on the derived sodium abundances. The resonance lines, 5889.951 Å and 5895.924 Å, are especially sensitive. Very limited NLTE abundances are available in the literature. The apparent large dispersion at [Fe/H] ¡ could probably be significantly reduced if NLTE corrections were applied. We used the INSPECT database666http://www.inspect-stars.com/ to compute corrections for our stars to investigate if the [Na/Fe] difference between Fornax and Carina would remain (see Table 9). The correction depends both on the measured LTE sodium abundances as well as the metallicity and the stellar atmospheric parameters. The corrections for Carina were [Na/Fe] and for Fornax [Na/Fe]. However, the Carina NLTE abundances are still higher, with [Na/Fe], and for Fornax [Na/Fe].
The interpretation of this Na-enhancement is not straight forward. Norris et al. (2013) found that CEMP-no stars were likely to be enhanced in Na, however, the Carina stars are depleted in carbon, see the previous Section, and Fig. 7. The origin of this Na enhancement, compared to other dSph galaxies, remains therefore unclear. We note however that there are two stars from Venn et al. (2012) with lower Na, so it seems that there a significant scatter of Na in Carina, similar to the Milky Way.
@@
The chemical abundances of the -elements Mg, and Ca, together with Ti, are shown in Fig. 9. All three elements show a very similar trend of enhanced at , in all our literature samples as well as in Hendricks et al. (2014a) who focuses on determining the position of the -knee in Fornax. Our results are in perfect agreement with the Milky Way, and other dwarf galaxies, both the dSphs (e.g. Tafelmeyer et al., 2010; Lucchesi et al., 2020; Theler et al., 2020) and UFDs (e.g. Simon, 2019).
The Carina dSph has a complex star formation history (e.g. Tolstoy et al., 2009; de Boer et al., 2014), and the scatter of [Mg/Fe] at intermediate metallicities, , is quite large, dex. However, the scatter seems to be reduced at , where all stars are consistent with the same [Mg/Fe] value, dex. It is therefore possible, that these low metallicities probe only the first star formation burst in Carina.
@@ 4.5 Iron-peak elements
Figure 10 presents Sc, Cr, Mn, Co, Ni, and Zn trends with Fe in Carina and Fornax, compared with the Milky Way halo and disc populations. The element Sc could only be determined in our Fornax sample, which had high-resolution UVES spectra. The production of Sc is dominated by core-collapse SNe (ccSN; Woosley et al. 2002; Battistini & Bensby 2015) and therefore, as expected, [Sc/Fe] is at the same level as the -elements seen in Fig. 9.
For both Fornax and Carina, Cr and Mn closely follow the Milky Way trends at as derived from 1D LTE methods. NLTE calculations for the neutral species of these elements show an over-ionization, leading to weakened lines and positive NLTE abundance corrections (Bergemann et al., 2019). Therefore the increasing trends of [Cr/Fe] and [Mn/Fe] with [Fe/H] might be artificial. In the case of Co, our measurements are lower than the average trend in the Milky Way. Similarly, the Co abundances in the Sculptor dSph are low at (Skúladóttir et al., 2023b).
Nickel can be produced in ccSN, as well as in thermonuclear supernovae Type Ia (SNIa; e.g., Jerkstrand, 2018). Before SNIa start to dominate, , our sample definitely set the level at solar value, , for both Carina and Fornax, implying that the global production of nickel follows that of iron in core-collapse supernovae. This is in excellent agreement with the Sculptor dSph and UFDs (Skúladóttir et al., 2023b). However, subsolar [Ni/Fe] has been noted in the Fornax population at (Letarte et al., 2010; Lemasle et al., 2014), and in Carina at (Norris et al., 2017). This clearly corresponds to a stage of the galaxy chemical evolution when the ejecta of SNIa dominate the composition of the interstellar medium. Similar results are found in other dSph galaxies (Hill et al., 2019; Kirby et al., 2019; Theler et al., 2020), as well as accreted dwarf galaxies (e.g. Nissen & Schuster, 2010).
There is a known increase of [Zn/Fe] towards low [Fe/H] (e.g. Cayrel et al., 2004), see Fig. 10 (bottom right). The NLTE corrections of Zn (Takeda et al., 2005) are small in the metal-poor regime and do not fully explain this trend, and it remains poorly understood. We were only able to measure Zn for the star fnx_06_019 which falls on the Milky Way halo trend at , comparable to the level of the [/Fe] plateau. This was also observed in two Sextans EMP stars (Lucchesi et al., 2020), two stars in Ursa Minor (Cohen & Huang, 2010), and in Sculptor (Skúladóttir et al., 2017; Skúladóttir et al., 2023b), highlighting the role of ccSN in the production of zinc in the early stage of galaxy evolution.
@@ 4.6 Neutron-capture elements
Figure 11 shows our measurements of [Sr/Fe] and [Ba/Fe] with [Fe/H]. The Ba and Sr abundances of our sample stars are consistent with the large spread in the Milky Way. The abundance ratio of [Sr/Ba] with [Ba/H] is presented in Fig. 12. The [Sr/Ba] ratios of the Fornax stars, are compatible with those of the Milky Way. However, these ratios in the Carina dSph are somewhat lower. Mashonkina et al. (2017b) identified two populations in [Sr/Ba], one with similar trend as the Milky Way (and Fornax), increasing towards low [Ba/H], and the other of near constant [Sr/Ba], similar to the empirical r-process ratio. The Carina abundances do seem to fall in between those two populations.
@@
Only in the r-I star, fnx_06_019, were we able to measure [Eu/Fe]. Its [Ba/Eu] = value sits just above the pure r-process ratio, as is shown in Fig. 13. This star, therefore likely also shows some minor s-process contribution, which is also evident from its rather high , see Fig. 12. Not only is this star rich in Eu, but it is also enhanced in La, Nd, and Dy, as evident from the very strong lines shown in Fig. 15, and listed Table 4. However, we note that the light neutron-capture elements Sr, Y, and Zr, are not enhanced, with , , and . The elements, Y, Sr, and Zr have a similar origin, and their ratios are observed to be approximately constant in the Milky Way halo (François et al., 2007).
@@
This is the first r-I star identified in a dSph galaxy at such low [Fe/H]. The r-I and r-II stars discovered in dSph galaxies so far seem to cover the more metal-rich end of the metallicity distribution compared to halo r-I stars (e.g. Shetrone et al., 2001; Aoki et al., 2007a; Cohen & Huang, 2009; Reichert et al., 2020).
5 Discussion & Conclusions
@@
Here we follow-up some of the first EMP candidates in the Fornax and Carina dSphs to populate the yet uncovered metallicity range . It is now clear that regardless of the subsequent evolution of the local classical dwarf galaxies, which harbor very different star formation histories, the first generations of stars formed in very similar way in these systems. Almost all chemical elements follow the same trend with (low) metallicity, and match the known relations for our Galaxy.
@@
However, there are clear differences in the [C/Fe] abundances of dSph galaxies, compared to UFDs and the Milky Way. In Fig. 14 we compare the cumulative fraction of CEMP-no stars in dSph galaxies to that of the Milky Way, in a similar way as Ji et al. (2020) did for UFDs. We used all the available literature data, corrected for carbon depletion due stellar evolution (Placco et al., 2014)777https://vplacco.pythonanywhere.com/, that had no or minimal selection effect in regards to C-abundance. In particular, we were unable to include Hansen et al. (2023), as they specifically targeted CEMP stars, thus not providing information about the CEMP-no fraction. No CEMP-s stars were included, however we did include the star in Canes Venatici I, without Ba measurement, which Yoon et al. (2020) claimed to be a CEMP-no star, using the stars position in the [Fe/H]-A(C) diagram. We note that [C/Fe] in this star is dex higher than all other dSph stars found in the literature. From Fig. 14 it is evident that the CEMP-no fraction in dSphs is significantly lower than in the Milky Way, reaching fractions of at , based on 15 CEMP-no stars, out of 251 in total. At the lowest metallicities, , none of the 7 known stars in dSph galaxies are C-enhanced, while the CEMP-no fraction in the Milky Way at these metallicities is . This also implies a discrepancy in the CEMP-fraction of dSphs and UFDs, since the latter are in agreement with the Milky Way (e.g. Ji et al., 2020). Our findings therefore confirm what has previously been stated with more limited data by e.g. Starkenburg et al. (2013); Skúladóttir et al. (2015); Skúladóttir et al. (2023b); Lucchesi et al. (2020).
@@
We see differences in the Na abundances of Fornax and Carina, where all the four stars in the Carina dSph, are high in , see Fig. 8. This is puzzling, since high Na abundances are typically found in CEMP-no stars (e.g. Norris et al., 2013), and are predicted by their theoretical yields (e.g. Heger & Woosley, 2002), but the Carina stars all have low . Furthermore, this is typically not seen in other very metal-poor stars in dSph galaxies (Tafelmeyer et al., 2010; Jablonka et al., 2015; Skúladóttir et al., 2023b).
@@
Like in the Milky Way, at low , we find a large scatter of the neutron-capture elements in both the Carina and the Fornax dSph galaxies. We report the discovery of a Eu-rich, r-I star in Fornax: , and . This star, fnx_06_019, also shows an outstanding enrichment in La, Nd and Dy ([X/Fe]¿+0.5). This is the first such case for a C-normal star at such low metallicity () in a classical dSph galaxy. This points to a prolific r-process event, early in the history of the Fornax dSph galaxy, where r-process enhancements were previously only seen at higher [Fe/H] ¿ –2 (Letarte et al., 2010; Lemasle et al., 2014; Reichert et al., 2020, 2021).
@@
The present analysis contributes to the limited understanding of the earliest chemical enrichment in dSph galaxies. To further investigate the range of abundance patterns observed in different systems, it is clear that more observations are needed. Fortunately, the situation would improve drastically with upcoming spectroscopic surveys, such as WEAVE (Jin et al., 2023) and 4MOST (de Jong et al., 2019), in particular the dedicated dwarf galaxy survey 4DWARFS (Skúladóttir et al., 2023a). With large and homogeneous data sets we will be able to identify rare EMP stars, and better constrain the CEMP-no fraction in different galaxies. Furthermore, we will get detailed spatial information, which will allow us to trace individual events, causing r-process enrichment. Our work serves to highlight that it is fundamental to study metal-poor stars in galaxies of different sizes and star formation histories to get a complete picture of the galaxy formation and first chemical enrichment.
@@ Acknowledgements.
- Hansen et al. (2018) Hansen, C. J., El-Souri, M., Monaco, L., et al. 2018, ApJ, 855, 83
@@ Appendix A Additional tables
Ti II | 5336.786 | 1.58 | – | – | 54.0 3.3 | – | – | – | – |

### v2

@@ 4.2 Carbon
Figure 7 presents the carbon abundances measured in Fornax and Carina at [Fe/H] ¡ 2.5, compared to the literature data. The four Carina stars analyzed in this paper all present very low carbon levels, similar to the upper limits derived in the Carina sample of Venn et al. (2012). The Carina [C/Fe] distribution is located at the lower edge of the Milky Way abundances, making Carina noticeably C-poor (see Fig. 4). However, we note that two CEMP-no stars have been identified in Carina, first in Susmitha et al. (2017) and more recently in Hansen et al. (2023).
In Fornax, our two stars and the one star previously studied by Tafelmeyer et al. (2010) have similar carbon-normal abundances. On average the Fornax stars are somewhat higher in C compared to Carina, especially when corrections for evolutionary status are taken into account.
We note that none of our six new stars is enhanced in carbon, while in the Milky Way the typical fraction of CEMP-no stars is at (Placco et al., 2014). This adds to previous results, showing that dSph galaxies are poor in CEMP-no stars relative to the Milky Way and the UFDs (Starkenburg et al., 2013; Skúladóttir et al., 2015, 2021, 2023a; Jablonka et al., 2015; Simon et al., 2015; Kirby et al., 2015; Hansen et al., 2018a; Lucchesi et al., 2020), for a more detailed discussion on this see Section 5.
@@ 4.3 Light element: Sodium
The LTE abundances of sodium are presented in Fig. 8. The Milky Way has a very large scatter in [Na/Fe] at low [Fe/H] ¡ , and our new data points fall within this range. Fornax shows a near solar ; instead, all four stars in Carina have high . Since direct comparison between the Carina (X-Shooter) and Fornax (UVES) spectra is not meaningful, in Fig. 4 we plot the Na lines of the Carina stars together with an X-Shooter spectrum of a Na-poor star in Sculptor. This comparison clearly shows the Carina stars to be Na rich.
Non-LTE effects can play an important role in the derived sodium abundances. The resonance lines, 5889.951 Å and 5895.924 Å, are especially sensitive. Very limited NLTE abundances are available in the literature. The apparent large dispersion at [Fe/H] ¡ could probably be significantly reduced if NLTE corrections were applied. We used the INSPECT database555http://www.inspect-stars.com/ to compute corrections for our stars to investigate whether the [Na/Fe] difference between Fornax and Carina would remain (see Table 2). The correction depends both on the measured LTE sodium abundances and on the metallicity and the stellar atmospheric parameters. The corrections for Carina were [Na/Fe] and for Fornax [Na/Fe]. However, the Carina NLTE abundances are still higher, with [Na/Fe]; for Fornax [Na/Fe].
The interpretation of this Na-enhancement is not straightforward. Norris et al. (2013) found that CEMP-no stars were likely to be enhanced in Na; however, the Carina stars are depleted in carbon (see the previous section, and Fig. 7). The origin of this Na enhancement, compared to other dSph galaxies, therefore remains unclear. We note however that there are two stars from Venn et al. (2012) with lower Na, so it seems that there is a significant scatter of Na in Carina, similar to the Milky Way.
@@
The chemical abundances of the -elements Mg and Ca, together with Ti, are shown in Fig. 9. All three elements show a very similar trend of enhanced at in all our literature samples, and in Hendricks et al. (2014a) who focuses on determining the position of the -knee in Fornax. Our results are in perfect agreement with the Milky Way, and other dwarf galaxies, both the dSph galaxies (e.g., Tafelmeyer et al., 2010; Lucchesi et al., 2020; Theler et al., 2020) and the UFDs (e.g., Simon, 2019).
The Carina dSph galaxy has a complex star formation history (e.g., Tolstoy et al., 2009; de Boer et al., 2014), and the scatter of [Mg/Fe] at intermediate metallicities, , is quite large, dex. However, the scatter seems to be reduced at , where all stars are consistent with the same [Mg/Fe] value, dex. It is therefore possible that these low metallicities probe only the first star formation burst in Carina.
@@ 4.5 Iron-peak elements
Figure 10 presents Sc, Cr, Mn, Co, Ni, and Zn trends with Fe in Carina and Fornax, compared with the Milky Way halo and disk populations. The element Sc could be determined only in our Fornax sample, which had high-resolution UVES spectra. The production of Sc is dominated by core-collapse supernovae (ccSNe; Woosley et al. 2002; Battistini & Bensby 2015), and therefore, as expected, [Sc/Fe] is at the same level as the -elements seen in Fig. 9.
For both Fornax and Carina, Cr and Mn closely follow the Milky Way trends at , as derived from 1D LTE methods. NLTE calculations for the neutral species of these elements show an overionization, leading to weakened lines and positive NLTE abundance corrections (Bergemann et al., 2019). Therefore, the increasing trends of [Cr/Fe] and [Mn/Fe] with [Fe/H] might be artificial. In the case of Co, our measurements are lower than the average trend in the Milky Way. Similarly, the Co abundances in the Sculptor dSph galaxy are low at (Skúladóttir et al., 2023b).
Nickel can be produced in ccSNe, and in thermonuclear Type Ia supernovae (SNIa; e.g., Jerkstrand, 2018). Before SNIa start to dominate, , our sample definitely set the level at solar value, , for both Carina and Fornax, implying that the global production of nickel follows that of iron in core-collapse supernovae. This is in excellent agreement with the Sculptor dSph galaxy and UFDs (Skúladóttir et al., 2023b). However, subsolar [Ni/Fe] has been noted in the Fornax population at (Letarte et al., 2010; Lemasle et al., 2014), and in Carina at (Norris et al., 2017). This clearly corresponds to a stage of the galaxy chemical evolution when the ejecta of SNIa dominate the composition of the interstellar medium. Similar results are found in other dSph galaxies (Hill et al., 2019; Kirby et al., 2019; Theler et al., 2020), as well as accreted dwarf galaxies (e.g., Nissen & Schuster, 2010).
There is a known increase in [Zn/Fe] toward low [Fe/H] (e.g., Cayrel et al., 2004) (see Fig. 10, bottom right). The NLTE corrections of Zn (Takeda et al., 2005) are small in the metal-poor regime and do not fully explain this trend, and it remains poorly understood. We were only able to measure Zn for the star fnx_06_019, which falls on the Milky Way halo trend at , comparable to the level of the [/Fe] plateau. This was also observed in two Sextans EMP stars (Lucchesi et al., 2020), two stars in Ursa Minor (Cohen & Huang, 2010), and in Sculptor (Skúladóttir et al., 2017, 2023b), highlighting the role of ccSNe in the production of zinc in the early stage of galaxy evolution.
@@ 4.6 Neutron-capture elements
Figure 11 shows our measurements of [Sr/Fe] and [Ba/Fe] with [Fe/H]. The Ba and Sr abundances of our sample stars are consistent with the large spread in the Milky Way. The abundance ratio of [Sr/Ba] with [Ba/H] is presented in Fig. 12. The [Sr/Ba] ratios of the Fornax stars are compatible with those of the Milky Way. However, these ratios in the Carina dSph galaxy are somewhat lower. Mashonkina et al. (2017b) identified two populations in [Sr/Ba], one with a trend similar to that found in the Milky Way (and Fornax), increasing toward low [Ba/H], and the other of near constant [Sr/Ba], similar to the empirical r-process ratio. The Carina abundances seem to fall between these two populations.
@@
Only in the r-I star, fnx_06_019, were we able to measure [Eu/Fe]. Its [Ba/Eu] = value sits just above the pure r-process ratio, as is shown in Fig. 13. This star, therefore, likely also shows some minor s-process contribution, which is evident from its rather high (see Fig. 12). Not only is this star rich in Eu, but it is also enhanced in La, Nd, and Dy, as is evident from the very strong lines shown in Fig. 15 and listed Table 4. However, we note that the light neutron-capture elements, Sr, Y, and Zr, are not enhanced: , , . The elements Y, Sr, and Zr have a similar origin, and their ratios are observed to be approximately constant in the Milky Way halo (François et al., 2007).
@@
This is the first r-I star identified in a dSph galaxy at such low [Fe/H]. The r-I and r-II stars discovered in dSph galaxies to date seem to cover the more metal-rich end of the metallicity distribution compared to halo r-I stars (e.g., Shetrone et al., 2001; Aoki et al., 2007a; Cohen & Huang, 2009; Hansen et al., 2018b; Reichert et al., 2020).
5 Discussion and conclusions
@@
Here we follow-up some of the first EMP candidates in the Fornax and Carina dSph galaxies to populate the as-yet-uncovered metallicity range . It is now clear that regardless of the subsequent evolution of the local classical dwarf galaxies, which harbor very different star formation histories, the first generations of stars formed in very similar way in these systems. Almost all the chemical elements follow the same trend with (low) metallicity, and match the known relations for our Galaxy.
@@
However, there are clear differences in the [C/Fe] abundances of dSph galaxies, compared to UFDs and the Milky Way. In Fig. 14 we compare the cumulative fraction of CEMP-no stars in dSph galaxies to that of the Milky Way, as Ji et al. (2020) did for UFDs. We used all the available literature data, corrected for carbon depletion due to stellar evolution (Placco et al., 2014),666https://vplacco.pythonanywhere.com/ that had minimal or no selection effects in regards to C-abundance. In particular, we were unable to include Hansen et al. (2023), as they specifically targeted CEMP stars, thus not providing information about the CEMP-no fraction. No CEMP-s stars (i.e. CEMP stars presenting s-process enhancement, [Ba/Fe] ¿ 0) were included; however, we did include the star in Canes Venatici I, without a Ba measurement, which Yoon et al. (2020) claimed was a CEMP-no star, using the stars position in the [Fe/H]-A(C) diagram. We note that [C/Fe] in this star is dex higher than all other dSph stars found in the literature. From Fig. 14 it is evident that the CEMP-no fraction in dSph galaxies is significantly lower than in the Milky Way, reaching fractions of at , based on 15 CEMP-no stars, out of 251 in total. At the lowest metallicities, , none of the seven known stars in dSph galaxies is C-enhanced, while the CEMP-no fraction in the Milky Way at these metallicities is . This also implies a discrepancy in the CEMP-no fraction of dSph galaxies and UFDs, since the latter are in agreement with the Milky Way fraction (e.g., Ji et al., 2020). Our findings therefore confirm what has previously been stated with more limited data (e.g., Starkenburg et al. 2013; Skúladóttir et al. 2015, 2023b; Lucchesi et al. 2020).
@@
We see differences in the Na abundances of Fornax and Carina, where all four stars in the Carina dSph galaxy are high in (see Fig. 8). This is puzzling since high Na abundances are typically found in CEMP-no stars (e.g., Norris et al., 2013), and are predicted by their theoretical yields (e.g., Heger & Woosley, 2002), but the Carina stars all have low . Furthermore, this is typically not seen in other very metal-poor stars in dSph galaxies (Tafelmeyer et al., 2010; Jablonka et al., 2015; Skúladóttir et al., 2023b).
@@
As we do in the Milky Way, at low we find a large scatter of the neutron-capture elements in both the Carina and the Fornax dSph galaxies. We report the discovery of a Eu-rich, r-I star in Fornax: and . This star, fnx_06_019, also shows an outstanding enrichment in La, Nd, and Dy ([X/Fe]¿+0.5). This is the first such case for a C-normal star at such low metallicity () in a classical dSph galaxy. This points to a prolific r-process event early in the history of the Fornax dSph galaxy, where r-process enhancements were previously only seen at higher [Fe/H] ¿ –2 (Letarte et al., 2010; Lemasle et al., 2014; Reichert et al., 2020, 2021).
@@
The present analysis contributes to the limited understanding of the earliest chemical enrichment in dSph galaxies. To further investigate the range of abundance patterns observed in different systems, it is clear that more observations are needed. Fortunately, the situation will improve drastically with upcoming spectroscopic surveys, such as WEAVE (Jin et al., 2023) and 4MOST (de Jong et al., 2019), in particular the dedicated dwarf galaxy survey 4DWARFS (Skúladóttir et al., 2023a). With large and homogeneous data sets we will be able to identify rare EMP stars, and better constrain the CEMP-no fraction in different galaxies. Furthermore, we will obtain detailed spatial information that will allow us to trace the individual events causing r-process enrichment. Our work serves to highlight that it is fundamental to study metal-poor stars in galaxies of different sizes and star formation histories to get a complete picture of the galaxy formation and first chemical enrichment.
@@ Acknowledgements.
- Hansen et al. (2018a) Hansen, C. J., El-Souri, M., Monaco, L., et al. 2018a, ApJ, 855, 83
- Hansen et al. (2018b) Hansen, C. J., El-Souri, M., Monaco, L., et al. 2018b, ApJ, 855, 83
@@ Appendix A Additional tables
| Ti II | 5336.786 | 1.58 | – | – | 54.0 3.3 | – | – | – | – | ||
| Fe I | Fe II | C** | O I | Na I | Mg I | Al I | Si I | Ca I | Sc II | Ti I | Ti II | Cr I | Mn I | Co I | Ni I | Cu I | Zn I | Sr II | Y II | Zr II | Ba II | La II | Pr II | Nd II | Eu II | Dy II | ||
| log(X)⊙ | 7.50 | 7.50 | 8.43 | 8.69 | 6.24 | 7.60 | 6.45 | 7.51 | 6.34 | 3.15 | 4.95 | 4.95 | 5.64 | 5.43 | 4.99 | 6.22 | 4.19 | 4.56 | 2.87 | 2.21 | 2.58 | 2.18 | 1.10 | 0.72 | 1.42 | 0.52 | 1.10 | |
| fnx_06_019 | ||||||||||||||||||||||||||||
| No. lines* | 67 | 2 | 1 | 1 | 2 | 3 | 1 | 1 | 6 | 6 | 9 | 7 | 4 | 3 | 1 | 3 | 1 | 1 | 1 | 5 | 1 | 4 | 4 | 1 | 3 | 2 | 3 | |
| log(X) | ¡ | |||||||||||||||||||||||||||
| [X/H] | ¡ | |||||||||||||||||||||||||||
| [X/Fe] | ¡ | |||||||||||||||||||||||||||
| Error | ||||||||||||||||||||||||||||
| fnx0579x–1 | ||||||||||||||||||||||||||||
| No. lines* | 76 | 4 | 1 | 1 | 2 | 4 | 1 | 11 | 5 | 10 | 8 | 5 | 3 | 1 | 4 | 1 | 1 | 2 | 4 | |||||||||
| log(X) | ||||||||||||||||||||||||||||
| [X/H] | ||||||||||||||||||||||||||||
| [X/Fe] | ||||||||||||||||||||||||||||
| Error | ||||||||||||||||||||||||||||
| car1_t174 | ||||||||||||||||||||||||||||
| No. lines* | 10 | 2 | 1 | 1 | 2 | 1 | 1 | 1 | 1 | 2 | ||||||||||||||||||
| log(X) | ||||||||||||||||||||||||||||
| [X/H] | ||||||||||||||||||||||||||||
| [X/Fe] | ||||||||||||||||||||||||||||
| Error | ||||||||||||||||||||||||||||
| car1_t194 | ||||||||||||||||||||||||||||
| No. lines* | 24 | 2 | 1 | 2 | 2 | 2 | 3 | 3 | 1 | 1 | 1 | 2 | ||||||||||||||||
| log(X) | ||||||||||||||||||||||||||||
| [X/H] | ||||||||||||||||||||||||||||
| [X/Fe] | ||||||||||||||||||||||||||||
| Error | ||||||||||||||||||||||||||||
| car1_t200 | ||||||||||||||||||||||||||||
| No. lines* | 12 | 1 | 2 | 2 | 2 | 1 | 1 | 1 | 2 | |||||||||||||||||||
| log(X) | ||||||||||||||||||||||||||||
| [X/H] | ||||||||||||||||||||||||||||
| [X/Fe] | ||||||||||||||||||||||||||||
| Error | ||||||||||||||||||||||||||||
| LG04c_0008 | ||||||||||||||||||||||||||||
| No. lines* | 35 | 6 | 1 | 2 | 2 | 1 | 2 | 1 | 1 | 2 | 1 | 1 | 2 | |||||||||||||||
| log(X) | ||||||||||||||||||||||||||||
| [X/H] | ||||||||||||||||||||||||||||
| [X/Fe] | ||||||||||||||||||||||||||||
| Error |